OpenStudy (anonymous):

3. Evaluate: 20 + 5 ÷ 5 - 3 (2 pts) 4. Simplify the expression: 2x - 3y + y - 3x (2pts) 10. Solve for x: -3(3x - 2) = -21 (2pts) 26. Find the slope between the 2 points. (2,5) and (4, 7) (2pts) 35. Solve for x. 2x - 5 > 3x + 4. (2 pts)

OpenStudy (anonymous):

3. 18

OpenStudy (hoblos):

3. 20 + 5÷5 -3 = 20 +1 -3 = 18 4. 2x - 3y + y - 3x= -x -2y 10. -3(3x - 2) = -21 -9x +6 = -21 9x= 27 x=3 26. slope = (7-5)/(4-2) = 1 35. 2x - 5 > 3x + 4 x<-9

OpenStudy (anonymous):

4. -x - 2 y

OpenStudy (anonymous):

aw! hoblos have already done it.